# Randomization Tests for Conditional Group Symmetry

### v1

@@ 5 Randomization tests for conditional group symmetry
The characterizations of conditional symmetry in the previous section suggest a simple way to test for conditional symmetry, in the form of a conditional randomization test (CRT). The basic idea for this setting is as follows. With a set of observations $(X,Y)_{1:n}$, we generate $B$ randomization sets $(X,Y)_{1:n}^{(b)}$ as
$(X,Y)_{i}^{(b)}\coloneq(\mathsf{\Gamma}_{i}^{(b)}\rho(X_{i}),\mathsf{\Gamma}_{i}^{(b)}\mathsf{\tilde{Y}}_{i}^{(b)})\;,\quad b=1,\dotsc,B.$
Proposition 4.2 indicates that the randomization might be implemented in three distinct ways, by conditioning on different parts of the observed data:
@@
    Condition on $\rho(X)_{1:n}$ and $\tilde{Y}_{1:n}$, with $\tilde{Y}_{i}=\tilde{\tau}^{-1}_{i}Y_{i}$, and where if necessary $\tilde{\tau}_{i}$ is sampled from the inversion kernel and fixed. Set $\mathsf{\tilde{Y}}_{i}^{(b)}=\tilde{Y}_{i}$ and sample $\mathsf{\Gamma}_{i}^{(b)}\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
@@
    Condition on $\rho(X)_{1:n}$ and $\tilde{\tau}_{1:n}$, where if necessary the latter are sampled from the inversion kernel and fixed. Set $\mathsf{\Gamma}_{i}^{(b)}=\tilde{\tau}_{i}$ and sample $\mathsf{\tilde{Y}}_{i}^{(b)}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
@@
    Condition on $\rho(X)_{1:n}$. Sample both $\mathsf{\Gamma}_{i}^{(b)}\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$ and $\mathsf{\tilde{Y}}_{i}^{(b)}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
Each of the procedures yields a different test; we differentiate between them by indicating what statistic of $(X,Y)_{1:n}$ is conditioned on, denoted by $S_{1:n}\coloneq S(X,Y)_{1:n}$. For example, in R1, $S_{i}=(\rho(X_{i}),\tilde{Y}_{i})$. Carrying out the appropriate randomization for each $i=1,\dotsc,n$ and repeating $B$ times, the CRT with test statistic $T$ estimates a conditional $p$-value as
@@
It is straightforward to show via Proposition 4.2 and its corollaries that the resulting CRT based on $B$ randomization sets is a level-$\alpha$ test. Moreover, as $B\to\infty$, the CRT produces a consistent conditional $p$-value, in the following sense. Under the null hypothesis that $P_{Y|X}$ is $\mathbf{G}$-equivariant, if $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$, then for $P$-almost every sequence $s_{1:n}$,
$p_{B}\;\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{\to}}\;\mathbb{P}\left(T\left((X^{\prime},Y^{\prime})_{1:n}\right))\geq T\left((X,Y)_{1:n}\right)\mid S^{\prime}_{1:n}=S_{1:n}=s_{1:n}\right)\;.$ (16)
Proposition 5.1 formalizes the result generally in terms of a test based on Proposition 4.2; the same result holds for a test based on Corollary 4.3. The proof is straightforward and provided in Section B.3 for completeness. As with the test of distributional invariance in Section 2.2, conditioning on $S_{1:n}$ means that $S_{1:n}$ can be used as inputs to the test statistic.
@@
Assume that $P_{Y|X}$ is $\mathbf{G}$-equivariant. For data $(X,Y)_{1:n}$, generate $B$ randomized versions $(X,Y)^{(b)}_{1:n}$ using any of the randomization procedures R1R3. Assume that $\mathbb{P}(T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})}))=0$ for $b\neq b^{\prime}$. For arbitrary fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, the statistic $p_{B}$ defined in (15) satisfies
[…]
@@ 5.1 Approximate conditional randomization
In Appendix C, we describe certain settings that allow for exact sampling from $P_{\Gamma|[X]}$. One particular case of interest described there is when $\mathbf{G}$ acts transitively on $\mathbf{X}$, in which case sampling from $P_{\Gamma|[X]}$ amounts to permuting the $\tilde{\tau}_{i}$’s. If $\mathbf{G}$ induces a finite number of orbits in $\mathbf{X}$, then an orbit-stratified permutation can be used.
@@
In general, we expect that sampling from approximations of $P_{\Gamma|[X]}$ and/or $P_{\tilde{Y}|[X]}$ will be the more common approach in practice. To that end, let $(X,Y)_{i,m}^{(b)}$ be a randomization of $(X,Y)_{i}$ generated as in Proposition 5.1, except that $\mathsf{\Gamma}_{i}^{(b)}$ and/or $\mathsf{\tilde{Y}}_{i}^{(b)}$ is sampled from an approximation to the appropriate conditional distribution. For example, this might be a sample from a kernel density estimate constructed from an independent dataset of $m$ observations. Under the null hypothesis, if as $m\to\infty$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)$ conditioned on $\rho(X_{i})$, then the approximate CRT will be asymptotically level-$\alpha$.
Proposition 5.2.
Assume $(X,Y)\sim P_{X,Y}$, with $P_{Y|X}$ $\mathbf{G}$-equivariant. Let $T$ be a continuous $\mathbb{R}$-valued function, such that the distribution of $T((X,Y)_{1:n})$ is atomless. Suppose that $(X,Y)_{i,m}^{(b)}$ are sampled conditionally on $S(X,Y)_{i,m}^{(b)}=S(X,Y)_{i}$, and used as the randomization sets to generate CRT statistics as in (15), denoted $p_{B,m}$. If, for each $i\leq n$, conditioned on $S(X,Y)_{i}$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)_{i}$ as $m\to\infty$, and $\mathbb{P}(T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})}))=0$ for $b\neq b^{\prime}$, then $p_{B,m}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}p_{B}$. Therefore, for fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$,
[…]
The proof can be found in Section B.3. We note that the sequence of approximations do not necessarily require an independent dataset to estimate the conditional distributions, but can depend on $(X,Y)_{1:n}$ through $S_{1:n}$. In that case, the limit in Proposition 5.2 should be taken as $n\to\infty$.
Any asymptotically correct method for approximating the required conditional distributions can be used for the conditional randomization procedure to yield an asymptotically level-$\alpha$ test. Here, we describe a kernel estimation approach for sampling from approximations of the conditional distributions required for R1, R2 and R3.
@@
Let $M(X)$ be any maximal invariant to condition on, such as $\rho(X)$, and let $k_{h_{m}}\colon\mathbf{M}\times\mathbf{M}\to\mathbb{R}_{+}$ be a kernel function defined on $\mathbf{M}$. Assume that $k_{h_{m}}$ is bounded above by $\kappa<\infty$, with $k_{h_{m}}(z,z)=\kappa$ for all $z\in\mathbf{M}$. Using a (possibly independent) dataset of size $m$ for estimation, assume that the bandwidth parameter $h_{m}$ decreases to 00 as $m\to\infty$, slowly enough to ensure asymptotically consistent density estimation. To generate a randomization conditioned on $M(X)$, we first assign to each index $i\in\{1,\dotsc,m\}$ a probability proportional to $k_{h_{m}}(M(X),M(X_{i}))$. To sample from an approximation of $P_{\tilde{Y}|[X]}$ for R2 and R3, we then sample an index $\ell$ from $\{1,\dotsc,m\}$ with those probabilities, and then take $\mathsf{\tilde{Y}}_{i}=\tilde{Y}_{\ell}$. Similarly, to sample from an approximation of $P_{\Gamma|[X]}$ for R1 and R3, we sample an index $j$ from $\{1,\dotsc,m\}$ (independent of $\ell$) with those probabilities, and then take $\mathsf{\Gamma}_{i}=\tilde{\tau}$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X_{j})$. It can be seen that as $m\to\infty$, the distributions from which this procedure samples approach the target conditional distributions. The resulting conditional randomization procedure (without an independent dataset) used in a test based on Properties E1, E2 and E3 is summarized in Algorithm 2. For a test based on Properties E4, E5 and E6, only the second component of the output pair is retained.
@@
5:     sample $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{j})$
@@
7:     sample $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X)$
@@
9:set $\mathsf{\Gamma}\coloneq\tilde{\tau}$
@@
12:else
[…]
14:end if
15:return $(\mathsf{\Gamma}\rho(X),\mathsf{\Gamma}\mathsf{\tilde{Y}})$
In our experiments, we estimate the conditional distributions using the same data used for testing. This procedure remains valid because we base the tests on Property E4, using only $\tilde{Y}_{1:n}$ and $Y_{1:n}$; the part of the observed data used for estimating conditional distributions, $\tilde{\tau}_{1:n}$, is conditionally independent of those variables under the null hypothesis. When the maximal invariant is continuous, we find that is it sufficient to use a Gaussian kernel with a diagonal covariance matrix given by Silverman’s rule-of-thumb (Silverman, 1986, Ch. 4).
@@ 5.2 Baseline: two-sample permutation tests
6 Kernel hypothesis tests
In the experiments reported in Section 7, we test for conditional symmetry using a kernel-based test statistic on observations denoted by $Z_{1:n}$. For tests based on Properties E1, E2 and E3, $Z_{i}=(X,Y)_{i}$, with a kernel defined on the product space $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$. For tests based on Properties E4, E5 and E6, $Z_{i}=Y_{i}$, with a kernel defined on $\mathbf{Z}=\mathbf{Y}$. We provide a brief overview of kernel-based hypothesis testing, and other related ideas in this section. See Christmann and Steinwart (2008) for a thorough treatment of the theory of kernels, and Muandet et al. (2017) for a review of their uses in statistics and machine learning.
Let $k_{\sigma}\colon\mathbf{Z}\times\mathbf{Z}\rightarrow\mathbb{R}$ be a positive semidefinite kernel with parameter $\sigma$, corresponding to the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$. If the kernel is measurable and bounded, then the kernel mean embedding (KME) of a probabililty measure $P$ on $\mathbf{Z}$ is the unique element $\mu_{P}({\,\vbox{\hbox{\tiny$\bullet$}}\,})\coloneqq\int_{\mathbf{Z}}k_{\sigma}(z,{\,\vbox{\hbox{\tiny$\bullet$}}\,})P(dz)\in\mathcal{H}$ such that $\mathbb{E}_{P}[f(Z)]=\langle f,\mu_{P}\rangle_{\mathcal{H}}$ for all $f\in\mathcal{H}$ (Gretton et al., 2012). It follows that $\langle\mu_{P_{1}},\mu_{P_{2}}\rangle=\int_{\mathbf{Z}\times\mathbf{Z}}k_{\sigma}(z,z^{\prime})P_{1}(dz)P_{2}(dz^{\prime})$. A kernel is said to be characteristic if the map $P\mapsto\mu_{P}$ from $\mathcal{P}(\mathbf{Z})$ into $\mathcal{H}$ is injective, meaning that the embedding for each probability measure is unique (Sriperumbudur et al., 2010). Kernel-based hypothesis tests compare distributions through their KMEs (Gretton et al., 2007). The test statistic in kernel-based tests is an estimator of the squared maximum mean discrepancy (MMD),
$\mathrm{MMD}_{\sigma}^{2}(P_{1},P_{2})\coloneqq\left\|\mu_{P_{1}}-\mu_{P_{2}}\right\|_{\mathcal{H}}^{2}$
If the kernel $k_{\sigma}$ is characteristic, $\mathrm{MMD}_{\sigma}^{2}$ is a metric on $\mathcal{P}(\mathbf{Z})$ (Gretton et al., 2012); hence $\mathrm{MMD}_{\sigma}(P,Q)=0$ if and only if $P=Q$. If a non-characteristic kernel is used instead, $\mathrm{MMD}_{\sigma}^{2}$ would be unable to separate some distinct elements of $\mathcal{P}(\mathbf{Z})$, which would not affect the level of the test, but may reduce the power under some alternatives.

### v2

@@ 4 Randomization tests for conditional group symmetry
The distributional identities in Section 3.2 provide the foundation for tests of conditional symmetry, in the form of conditional randomization tests (CRTs). The basic idea is as follows. With a set of observations $(X,Y)_{1:n}$, generate $B$ randomization sets $(X,Y)_{1:n}^{(b)}$ as
$(X,Y)_{i}^{(b)}\coloneq(\mathsf{\Gamma}_{i}^{(b)}\rho(X_{i}),\mathsf{\Gamma}_{i}^{(b)}\mathsf{\tilde{Y}}_{i}^{(b)})\;,\quad b=1,\dotsc,B,\ \ i=1,\dotsc,n.$
The distributional identities in Proposition C.1 indicate that the randomization can be implemented in three distinct ways by conditioning on observed data differently:
@@
    Condition on $\rho(X)_{1:n}$ and $\tilde{Y}_{1:n}$, with $\tilde{Y}_{i}=\Psi^{-1}_{i}Y_{i}$, and where $\Psi_{i}$ is a representative inversion if one exists or is otherwise sampled from the inversion kernel and fixed. Set $\mathsf{\tilde{Y}}_{i}^{(b)}=\tilde{Y}_{i}$ and sample $\mathsf{\Gamma}_{i}^{(b)}|\rho(X_{i})\sim P_{\Gamma|R}$.
@@
    Condition on $\rho(X)_{1:n}$ and $\Psi_{1:n}$, where the latter is a representative inversion if one exists or is otherwise sampled from the inversion kernel and fixed. Set $\mathsf{\Gamma}_{i}^{(b)}=\Psi_{i}$ and sample $\mathsf{\tilde{Y}}_{i}^{(b)}|\rho(X_{i})\sim P_{\tilde{Y}|R}$.
@@
    Condition on $\rho(X)_{1:n}$. Sample both $\mathsf{\Gamma}_{i}^{(b)}|\rho(X_{i})\sim P_{\Gamma|R}$ and $\mathsf{\tilde{Y}}_{i}^{(b)}|\rho(X_{i})\sim P_{\tilde{Y}|R}$.
In Section D.1, we show that in settings where a representative inversion $\psi^{\circ}$} can be defined despite the group action not being free, randomizing to $\Psi$ is not necessary in order for the tests to have the appropriate level.
Each of the randomization procedures above yields a different test; we differentiate between them by indicating what statistic of $(X,Y)_{1:n}$ is conditioned on, denoted $S_{1:n}\coloneq S(X,Y)_{1:n}$. For example, in R1, $S_{i}=(\rho(X_{i}),\tilde{Y}_{i})$. Carrying out the appropriate randomization for each $i=1,\dotsc,n$ and repeating $B$ times, the CRT with test statistic $T$ estimates a conditional $p$-value as
@@
It is straightforward to show via Proposition C.1 and its corollaries that the resulting CRT based on $B$ randomization sets is a level-$\alpha$ test. Moreover, as $B\to\infty$, the CRT produces a consistent conditional $p$-value, in the following sense. Under the null hypothesis that $P_{Y|X}$ is $\mathbf{G}$-equivariant, if $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$, then for $P$-almost every sequence $s_{1:n}$, as $B\to\infty$,
$p_{B}\;\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{\to}}\;\mathbb{P}\left\{T\left((X^{\prime},Y^{\prime})_{1:n}\right))\geq T\left((X,Y)_{1:n}\right)\mid S^{\prime}_{1:n}=S_{1:n}=s_{1:n}\right\}\;.$ (12)
Proposition 4.1 formalizes the result generally in terms of a test based on the identity (9) (from Proposition C.1); the same holds for a test based on the identity (10) (from Corollary C.2). The proof is straightforward and provided in Section C.3 for completeness. As with the test of distributional invariance in Section 2.2, conditioning on $S_{1:n}$ means that $S_{1:n}$ can be used as input to the test statistic.
@@
Assume that $P_{Y|X}$ is $\mathbf{G}$-equivariant. For data $(X,Y)_{1:n}$, generate $B$ randomized versions $(X,Y)^{(b)}_{1:n}$ using any of the randomization procedures R1R3. Assume that $\mathbb{P}\{T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})})\}=0$ for $b\neq b^{\prime}$. For arbitrary fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, the statistic $p_{B}$ defined in (11) satisfies
$\mathbb{P}\left(p_{B}\leq\alpha\mid S_{1:n}\right)=\frac{\lfloor\alpha(B+1)\rfloor}{B+1}\leq\alpha\;,\quad P\text{-a.s.}$ (13)
@@
As mentioned in Section 2.2, one sets $B$ before conducting the test. However, there are approaches that allow $B$ to be chosen adaptively, which may have benefits in terms of improving computational efficiency while retaining test validity; see (Stoepker and Castro, 2024) for a recent survey. Our tests may also benefit from these approaches, though because simulating samples from the relevant conditional distributions is not overly expensive, we do not pursue that here.
@@ 4.1 Approximate conditional randomization
In Section D.2, we describe certain settings that allow for exact sampling from $P_{\Gamma|R}$. One particular case of interest described there is when $\mathbf{G}$ acts transitively on $\mathbf{X}$, in which case sampling from $P_{\Gamma|R}$ amounts to permuting the $\Psi_{i}$’s. If $\mathbf{G}$ induces a finite number of orbits in $\mathbf{X}$, then an orbit-stratified permutation can be used. In general, we expect that sampling from approximations of $P_{\Gamma|R}$ and/or $P_{\tilde{Y}|R}$ will be the more common approach in practice.
@@
To that end, let $(X,Y)_{i,m}^{(b)}$ be a randomization of $(X,Y)_{i}$ generated as in Proposition 4.1, except that $\mathsf{\Gamma}_{i}^{(b)}$ and/or $\mathsf{\tilde{Y}}_{i}^{(b)}$ is sampled from an approximation to the appropriate conditional distribution. For example, this might be a sample from a kernel density estimate constructed from an independent dataset of $m$ observations. Under the null hypothesis, if $T$ is a continuous function and if as $m\to\infty$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)$ conditioned on $\rho(X_{i})$, then the approximate CRT will be asymptotically level-$\alpha$. We state and prove this formally in Section C.3. We note that the sequence of approximations does not necessarily require an independent dataset to estimate the conditional distributions, but can depend on $S_{1:n}$. Hence, any asymptotically correct method for approximating the required conditional distributions can be used for the conditional randomization procedure to yield an asymptotically level-$\alpha$ test.
@@
Here, we describe a kernel estimation approach for sampling from approximations of the conditional distributions required for R1, R2 and R3. Let $M(X)$ be a continuous maximal invariant, such as $\rho(X)$ chosen so that $\rho$ is a continuous function, and let $k_{h_{m}}\colon\mathbf{M}\times\mathbf{M}\to\mathbb{R}_{+}$ be a kernel function defined on $\mathbf{M}$. Assume that $k_{h_{m}}$ is bounded above by $\nu<\infty$, with $k_{h_{m}}(z,z)=\nu$ for all $z\in\mathbf{M}$. Using a (possibly independent) dataset of size $m$ for estimation, assume that the bandwidth parameter $h_{m}$ decreases to 0 as $m\to\infty$, slowly enough to ensure asymptotically consistent density estimation. To generate a randomization conditioned on $M(X)$, we first assign to each index $i\in\{1,\dotsc,m\}$ a probability proportional to $k_{h_{m}}(M(X),M(X_{i}))$. To sample from an approximation of $P_{\tilde{Y}|R}$ for R2 and R3, we then sample an index $\ell$ from $\{1,\dotsc,m\}$ with those probabilities, and then take $\mathsf{\tilde{Y}}_{i}=\tilde{Y}_{\ell}$. Similarly, to sample from an approximation of $P_{\Gamma|R}$ for R1 and R3, we sample an index $j$ from $\{1,\dotsc,m\}$ (independent of $\ell$) with those probabilities, and then take $\mathsf{\Gamma}_{i}=\Psi$, where $\Psi|X_{j}\sim P_{\Psi|X}$. It can be seen that as $m\to\infty$, the distributions from which this procedure samples approach the target conditional distributions. The resulting conditional randomization procedure (without an independent dataset) used in a test based on Proposition C.1 is summarized in Algorithm 2. For a test based on Corollary C.2, only the second component of the output pair is retained.
@@
5:  sample $\Psi|X_{j}\sim P_{\Psi|X}$
@@
7:  sample $\Psi|X\sim P_{\Psi|X}$
@@
9:set $\mathsf{\Gamma}\coloneq\Psi$
@@
11:  set $\mathsf{\tilde{Y}}\coloneq\tilde{Y}_{\ell}$
12:else
13:  set $\mathsf{\tilde{Y}}\coloneq\tilde{Y}$
14:end if
@@
In our experiments, we estimate the conditional distributions using the same data used for testing. This procedure remains valid because we base the tests on Corollary C.2, using only $\tilde{Y}_{1:n}$; the part of the observed data used for estimating conditional distributions, $(\Psi_{1:n},M(X)_{1:n})$, is conditionally independent of those variables under the null hypothesis. With a continuous maximal invariant, we find that it is sufficient to use a Gaussian kernel with a diagonal covariance matrix given by Silverman’s rule of thumb (Silverman, 1986, Ch. 4). Our experiments show that with finite samples, using Silverman’s rule for approximate randomization tends to yield conservative tests with sizes less than $\alpha$. Such tests are likely to only achieve exact level-$\alpha$ as $m\to\infty$ or if exact randomization is possible in settings like those described in Section D.2. Using larger bandwidths can increase power, although it can also lead to inflated Type I error rates (see Section E.2).
@@ 4.2 Baseline: two-sample permutation tests
5 Kernel-based hypothesis tests for conditional group symmetry
In the experiments reported in Section 6, we test for conditional symmetry using a kernel-based test statistic of observations denoted for convenience by $Z_{1:n}$. For tests based on identity (9) (Proposition C.1), $Z_{i}=(X,Y)_{i}$, with a kernel defined on the product space $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$. For tests based on (10) (Corollary C.2), $Z_{i}=Y_{i}$, with a kernel defined on $\mathbf{Z}=\mathbf{Y}$. We provide a brief overview of kernel-based hypothesis testing and other related ideas in this section. See Christmann and Steinwart (2008) for a thorough treatment of the theory of kernels, and Muandet et al. (2017) for a review of their uses in statistics and machine learning.
Let $k_{\sigma}\colon\mathbf{Z}\times\mathbf{Z}\rightarrow\mathbb{R}$ be a positive semidefinite kernel with parameter $\sigma$, corresponding to the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$. If the kernel is measurable and bounded, then the kernel mean embedding (KME) of a probability measure $P$ on $\mathbf{Z}$ is the unique element $\mu_{P}({\,\vbox{\hbox{\tiny$\bullet$}}\,})\coloneqq\int_{\mathbf{Z}}k_{\sigma}(z,{\,\vbox{\hbox{\tiny$\bullet$}}\,})P(dz)\in\mathcal{H}$ such that $\mathbb{E}_{P}[f(Z)]=\langle f,\mu_{P}\rangle_{\mathcal{H}}$ for all $f\in\mathcal{H}$ (Gretton et al., 2012). A kernel is said to be characteristic if the map $P\mapsto\mu_{P}$ from $\mathcal{P}(\mathbf{Z})$ into $\mathcal{H}$ is injective, meaning that the embedding for each probability measure is unique (Sriperumbudur et al., 2010). KME-based hypothesis tests compare distributions through their KMEs (Gretton et al., 2007, 2012). The test statistic in KME-based tests is an estimator of the squared maximum mean discrepancy (MMD), $\mathrm{MMD}_{\sigma}^{2}(P_{1},P_{2})\coloneqq\left\|\mu_{P_{1}}-\mu_{P_{2}}\right\|_{\mathcal{H}}^{2}$. If the kernel $k_{\sigma}$ is characteristic, $\mathrm{MMD}_{\sigma}^{2}$ is a metric on $\mathcal{P}(\mathbf{Z})$; hence $\mathrm{MMD}_{\sigma}(P,Q)=0$ if and only if $P=Q$. If a non-characteristic kernel is used instead, $\mathrm{MMD}_{\sigma}^{2}$ may be unable to separate some distinct elements of $\mathcal{P}(\mathbf{Z})$, which would not affect the level of the test, but may reduce the power under some alternatives.